### v1

@@ 4.1 Datasets, metrics, and experimental settings
4.2 Hyperparameter Experiments of HAL
This subsection investigates the key hyperparameters that influence the scalability, diversity, and uncertainty of the proposed HAL module, while keeping all other training configurations identical to those described in Section 4.1. These include the AdamW optimizer, batch size, number of epochs, EMA decay rate of 0.99, and the DINOv2-Small encoder with the same classifier within the iterative loops under corresponding labeled budgets. All results are reported with OA and AA.
4.2.1 Effect of Mini-Batch Size (Scalability)
To evaluate the scalability of the HAL for sample selection, we vary the mini-batch size $S$ used in the HAL stage on the large-scale NWPU-RESISC45, while keeping the rest of HAL intact. Specifically, we test $S\in\{1000,\,3000,\,5000,\,$ 7000, 10000,7000,\,10000, 15000, 20000}15000,\,20000\}. The cluster multiplier $\lambda$ in Eq. (11) is fixed as 3.0. We report OA/AA and runtime to characterize the selection quality and computational cost. This experiment highlights HAL’s scalability achieved via mini-batch partitioning. Results are summarized in The results are summarized in Table 1, and the total runtime of spectral clustering for different unlabeled partition under the 8% labeled ratio is illustrated in Fig. 2.
Table 1: Evaluation of HAL’s scalability on NWPU-RESISC45 under different mini-batch sizes, reporting overall accuracy (OA), average accuracy (AA), and the runtime (seconds) required to perform spectral clustering on the entire unlabeled pool during each round. The best OA and AA at each labeling ratio (except 1%) are highlighted in bold. Here, “–” indicates that no results are reported because the runtime was excessively long, causing the sklearn spectral clustering algorithm to terminate automatically.
@@ 4.2.1 Effect of Mini-Batch Size (Scalability)
As shown in Table 1, overall, increasing the mini-batch size from 1000 to 5000 improves both OA and AA across all labeling ratios while maintaining moderate runtime. The mini-batch strategy, which randomly partitions the unlabeled pool into multiple subsets and performs spectral clustering within each subset, effectively preserves selection quality and diversity while greatly reducing computational complexity. This design allows HAL to maintain both global representativeness and local diversity among queried samples, making spectral clustering practical even for large-scale unlabeled datasets.
@@
4.2.2 Effect of Cluster Multiplier $\lambda$ (Diversity)
To investigate the impact of clustering granularity in the spectral clustering stage of HAL, we examine the cluster multiplier $\lambda$, which determines the number of spectral clusters in each mini-batch as $K_{b}=\lambda B_{b}$ in Eq. (11), where $B_{b}$ denotes the allocated query budget. By varying $\lambda$, we control the fineness of spectral clustering, thereby influencing the diversity of queried samples and the overall clustering quality. Experiments are conducted on the UCM and AID datasets under the HAL configuration with $\lambda\in\{1.0,\ 2.0,\ 3.0,\ 4.0\}$. For each $\lambda$, the complete HAL loop is executed, with OA and AA reported across all annotation rounds to evaluate how $\lambda$ affects spectral clustering behavior and sample diversity. Under different values of $\lambda$, the quantitative results are summarized in Table 2, while Fig. 3 visualizes the corresponding distributions of spectral cluster sizes based on 2% of the AID dataset.
Table 2: HAL’s performance under different values of $\lambda$ on UCM and AID datasets, reporting overall accuracy (OA) and average accuracy (AA). The best OA and AA at each labeling ratio (except 1%) are highlighted in bold.
@@ 4.2.2 Effect of Cluster Multiplier $\lambda$ (Diversity)
As shown in Table 2, varying the cluster multiplier $\lambda$ affects the granularity of spectral clustering and thereby the diversity of the queried samples. When $\lambda$ increases from 1.0 to 3.0, both OA and AA generally improve across UCM and AID, indicating that moderate cluster expansion enhances the representation of local structures and improves sample diversity. However, an excessively large $\lambda$ (e.g., 4.0) tends to over-fragment the clusters, as also reflected in Fig. 3, leading to unstable clustering and slight performance degradation.
@@
4.2.3 In-Cluster Selection Criterion (Uncertainty)
To examine the role of HAL’s uncertainty in guiding in-cluster sample selection, we compare three strategies on UCM and AID: (i) S-Random selection, which randomly picks one sample per cluster; (ii) S-Centroid selection, which chooses the sample closest to the cluster center, representing the most representative instance; and (iii) S-Uncertainty selection, which selects the sample with the highest uncertainty, measured by the gradient norm as described in Sec. 3.3.3, to focus on ambiguous or hard-to-classify regions. All other settings of HAL remain the same. This comparison allows us to assess whether incorporating uncertainty helps the model identify more informative samples beyond those merely representative of the cluster structure.
As shown in Table 3, the S-Centroid strategy already performs better than the S-Random strategy, as selecting samples closer to cluster centers yields more representative and stable supervision. However, our S-Uncertainty strategy further improves both OA and AA across all annotation ratios on UCM and AID in general. It achieves higher OA, indicating superior overall classification performance, and consistently higher AA, reflecting better class balance among the queried samples. This demonstrates that uncertainty-guided sampling enables HAL to focus on informative and ambiguous regions near decision boundaries while maintaining representative coverage, thereby improving both efficiency and fairness in sample selection.
Table 3: HAL’s performance of different sample selection strategies on UCM and AID datasets, reporting overall accuracy (OA) and average accuracy (AA). The best OA and AA at each labeling ratio (except 1%) are highlighted in bold.
@@ 4.2.3 In-Cluster Selection Criterion (Uncertainty)
Table 4: Ablation study of HSSAL on UCM and AID datasets, reporting overall accuracy (OA) and average accuracy (AA). The best OA and AA at each labeling ratio (except 1%) are highlighted in bold.
@@
4.3 Ablation study of HSSAL
@@
As shown in Table 4 and Fig. 4, the ablation results reveal that the individual contributions of SSL and AL differ notably between datasets of different scales. On the small-scale UCM dataset, introducing FixMatch (Random+FixMatch) for self-training on unlabeled data does not consistently improve performance compared with the supervised-only baseline (Random); in fact, it even degrades the OA under low labeling ratios, likely due to overfitting to noisy pseudo-labels caused by limited data diversity. In contrast, the HAL querying strategy shows clear gains over random selection, indicating its ability to identify more informative and representative samples. On the larger and more diverse AID dataset, the trend reverses: FixMatch provides significant improvements over random sampling, while the pure HAL variant contributes less prominently. This suggests that FixMatch benefits more from abundant unlabeled data, whereas HAL’s effectiveness is constrained when sample diversity is already high.
@@
Furthermore, the visualization results of unlabeled training samples under the 2% labeled setting based on the models trained by Random and Random+FixMatch are illustrated in Fig. 5, using the UMAP technique (mcinnes2018umap). Fig. 5 presents the true distribution of the Random strategy, while Figs. 5 and 5 visualize the unlabeled samples selected by Random and HAL, respectively. In addition, Fig. 5 displays the spectral clustering results with 140 clusters (for clarity, only 140 clusters are shown instead of 420 corresponding to $\lambda=3$), where the clustering process partitions the dataset into multiple locally coherent subsets, providing an effective foundation for HAL’s subsequent informative sample selection shown in Fig. 5. The comparison between Fig. 5 and Figs. 55 indicates that SSL can effectively enhance the discrimination of class boundaries. Moreover, by comparing Figs. 5 and 5, it can be observed that HAL tends to select more informative samples that are both diverse across the entire dataset and highly uncertain near the class boundaries.
4.4 Comparison experiments
@@
Table 5: Comparison results on the UCM dataset among HAL, various AL strategies, SSL methods, and their integrated variants (HSSAL). The table reports overall accuracy (OA) and average accuracy (AA) under different labeling ratios. The best OA and AA at each ratio (except 1%) and their average values are highlighted in bold.
@@
Table 6: Comparison results on the AID dataset among HAL, various AL strategies, SSL methods, and their integrated variants (HSSAL). The table reports overall accuracy (OA) and average accuracy (AA) under different labeling ratios. The best OA and AA at each ratio (except 1%) and their average values are highlighted in bold.
@@
Table 7: Comparison results on the NWPU-RESISC45 dataset among HAL, various AL strategies, SSL methods, and their integrated variants (HSSAL). The table reports overall accuracy (OA) and average accuracy (AA) under different labeling ratios. The best OA and AA at each ratio (except 1%) and their average values are highlighted in bold.
@@
  • Entropy (settles2009active): Queries samples with the highest predictive entropy, reflecting maximum model uncertainty.
  • Margin (scheffer2001active): Selects samples with the smallest difference between the top two class probabilities, focusing on ambiguous predictions near the decision boundary.
  • BALD (bald): Estimates both epistemic and aleatoric uncertainty by maximizing the mutual information between predictions and model parameters. We adopt Monte Carlo dropout with $T=50$ stochastic forward passes on the last layer to approximate the acquisition score.
  • CoreSet (coreset): Selects representative samples via $k$-center clustering in the feature space. The image features are the last-layer [CLS] tokens from the DINOv2-Small encoder, identical to those used in HAL.
  • BADGE (badge): Combines uncertainty and diversity by clustering gradient embeddings with $k$-means++. It operates on the same DINOv2 [CLS] features as CoreSet and HAL.
@@
  • FixMatch (sohn2020fixmatch): A representative weak-to-strong consistency method that generates pseudo-labels from weakly augmented images and enforces consistency on their strongly augmented counterparts when confidence exceeds a fixed threshold $\tau=0.95$.
  • FlexMatch (zhang2021flexmatch): Builds on FixMatch by introducing a class-wise adaptive threshold $\tau_{c}$ according to the learning progress of each class to reduce confirmation bias.
  • FreeMatch (wang2023freematch): Further improves FixMatch by dynamically adjusting both global and class-specific thresholds based on prediction history for adaptive pseudo-labeling.
  • SoftMatch (chen2023softmatch): Extends the FixMatch paradigm by replacing the hard threshold with Gaussian confidence weighting, enabling smoother exploitation of uncertain samples.
As shown in Tables 57, among all AL baselines, the proposed HAL consistently achieves the highest OA and AA across different labeling ratios. The close alignment between OA and AA indicates that HAL not only selects the most informative samples but also maintains class balance during the annotation process. This advantage mainly stems from its hierarchical sample selection mechanism, which ensures both representativeness and diversity in sample selection from the unlabeled pool.

### v2

@@ IV-A Datasets, metrics, and experimental settings
IV-B Hyperparameter Experiments of HAL
This subsection investigates the key hyperparameters that influence the scalability, diversity, and uncertainty of the proposed HAL module, while keeping all other training configurations identical to those described in Section IV-A. These include the AdamW optimizer, batch size, number of epochs, EMA decay rate of 0.99, and the DINOv2-Small encoder with the same classifier within the iterative loops under corresponding labeled budgets. All results are reported with OA and AA.
IV-B1 Effect of Mini-Batch Size (Scalability)
To evaluate the scalability of the HAL for sample selection, we vary the mini-batch size $S$ used in the HAL stage on the large-scale NWPU-RESISC45, while keeping the rest of HAL intact. Specifically, we test $S\in\{1000,\,3000,\,5000,\,$ 7000, 10000,7000,\,10000, 15000, 20000}15000,\,20000\}. The cluster multiplier $\lambda$ in Eq. (11) is fixed as 3.0. We report OA/AA and runtime to characterize the selection quality and computational cost. This experiment highlights HAL’s scalability achieved via mini-batch partitioning. Results are summarized in The results are summarized in Table I, and the total runtime of spectral clustering for different unlabeled partition under the 8% labeled ratio is illustrated in Fig. 2.
TABLE I: Evaluation of HAL’s scalability on NWPU-RESISC45 under different mini-batch sizes, reporting overall accuracy (OA), average accuracy (AA), and the runtime (seconds) required to perform spectral clustering on the entire unlabeled pool during each round. The best OA and AA at each labeling ratio (except 1%) are highlighted in bold. Here, “–” indicates that no results are reported because the runtime was excessively long, causing the sklearn spectral clustering algorithm to terminate automatically.
@@ IV-B1 Effect of Mini-Batch Size (Scalability)
As shown in Table I, overall, increasing the mini-batch size from 1000 to 5000 improves both OA and AA across all labeling ratios while maintaining moderate runtime. The mini-batch strategy, which randomly partitions the unlabeled pool into multiple subsets and performs spectral clustering within each subset, effectively preserves selection quality and diversity while greatly reducing computational complexity. This design allows HAL to maintain both global representativeness and local diversity among queried samples, making spectral clustering practical even for large-scale unlabeled datasets.
@@
IV-B2 Effect of Cluster Multiplier $\lambda$ (Diversity)
To investigate the impact of clustering granularity in the spectral clustering stage of HAL, we examine the cluster multiplier $\lambda$, which determines the number of spectral clusters in each mini-batch as $K_{b}=\lambda B_{b}$ in Eq. (11), where $B_{b}$ denotes the allocated query budget. By varying $\lambda$, we control the fineness of spectral clustering, thereby influencing the diversity of queried samples and the overall clustering quality. Experiments are conducted on the UCM and AID datasets under the HAL configuration with $\lambda\in\{1.0,\ 2.0,\ 3.0,\ 4.0\}$. For each $\lambda$, the complete HAL loop is executed, with OA and AA reported across all annotation rounds to evaluate how $\lambda$ affects spectral clustering behavior and sample diversity. Under different values of $\lambda$, the quantitative results are summarized in Table II, while Fig. 3 visualizes the corresponding distributions of spectral cluster sizes based on 2% of the AID dataset.
TABLE II: HAL’s performance under different values of $\lambda$ on UCM and AID datasets, reporting overall accuracy (OA) and average accuracy (AA). The best OA and AA at each labeling ratio (except 1%) are highlighted in bold.
@@ IV-B2 Effect of Cluster Multiplier $\lambda$ (Diversity)
As shown in Table II, varying the cluster multiplier $\lambda$ affects the granularity of spectral clustering and thereby the diversity of the queried samples. When $\lambda$ increases from 1.0 to 3.0, both OA and AA generally improve across UCM and AID, indicating that moderate cluster expansion enhances the representation of local structures and improves sample diversity. However, an excessively large $\lambda$ (e.g., 4.0) tends to over-fragment the clusters, as also reflected in Fig. 3, leading to unstable clustering and slight performance degradation.
@@
IV-B3 In-Cluster Selection Criterion (Uncertainty)
To examine the role of HAL’s uncertainty in guiding in-cluster sample selection, we compare three strategies on UCM and AID: (i) S-Random selection, which randomly picks one sample per cluster; (ii) S-Centroid selection, which chooses the sample closest to the cluster center, representing the most representative instance; and (iii) S-Uncertainty selection, which selects the sample with the highest uncertainty, measured by the gradient norm as described in Sec. III-C3, to focus on ambiguous or hard-to-classify regions. All other settings of HAL remain the same. This comparison allows us to assess whether incorporating uncertainty helps the model identify more informative samples beyond those merely representative of the cluster structure.
As shown in Table III, the S-Centroid strategy already performs better than the S-Random strategy, as selecting samples closer to cluster centers yields more representative and stable supervision. However, our S-Uncertainty strategy further improves both OA and AA across all annotation ratios on UCM and AID in general. It achieves higher OA, indicating superior overall classification performance, and consistently higher AA, reflecting better class balance among the queried samples. This demonstrates that uncertainty-guided sampling enables HAL to focus on informative and ambiguous regions near decision boundaries while maintaining representative coverage, thereby improving both efficiency and fairness in sample selection.
TABLE III: HAL’s performance of different sample selection strategies on UCM and AID datasets, reporting overall accuracy (OA) and average accuracy (AA). The best OA and AA at each labeling ratio (except 1%) are highlighted in bold.
@@ IV-B3 In-Cluster Selection Criterion (Uncertainty)
TABLE IV: Ablation study of HSSAL on UCM and AID datasets, reporting overall accuracy (OA) and average accuracy (AA). The best OA and AA at each labeling ratio (except 1%) are highlighted in bold.
@@
IV-C Ablation study of HSSAL
@@
As shown in Table IV and Fig. 4, the ablation results reveal that the individual contributions of SSL and AL differ notably between datasets of different scales. On the small-scale UCM dataset, introducing FixMatch (Random+FixMatch) for self-training on unlabeled data does not consistently improve performance compared with the supervised-only baseline (Random); in fact, it even degrades the OA under low labeling ratios, likely due to overfitting to noisy pseudo-labels caused by limited data diversity. In contrast, the HAL querying strategy shows clear gains over random selection, indicating its ability to identify more informative and representative samples. On the larger and more diverse AID dataset, the trend reverses: FixMatch provides significant improvements over random sampling, while the pure HAL variant contributes less prominently. This suggests that FixMatch benefits more from abundant unlabeled data, whereas HAL’s effectiveness is constrained when sample diversity is already high.
@@
Furthermore, the visualization results of unlabeled training samples under the 2% labeled setting based on the models trained by Random and Random+FixMatch are illustrated in Fig. 5, using the UMAP technique[65]. Fig. 5a presents the true distribution of the Random strategy, while Figs. 5c and 5d visualize the unlabeled samples selected by Random and HAL, respectively. In addition, Fig. 5b displays the spectral clustering results with 140 clusters (for clarity, only 140 clusters are shown instead of 420 corresponding to $\lambda=3$), where the clustering process partitions the dataset into multiple locally coherent subsets, providing an effective foundation for HAL’s subsequent informative sample selection shown in Fig. 5d. The comparison between Fig. 5a and Figs. 5c5d indicates that SSL can effectively enhance the discrimination of class boundaries. Moreover, by comparing Figs. 5c and 5d, it can be observed that HAL tends to select more informative samples that are both diverse across the entire dataset and highly uncertain near the class boundaries.
IV-D Comparison experiments
@@
TABLE V: Comparison results on the UCM dataset among HAL, various AL strategies, SSL methods, and their integrated variants (HSSAL). The table reports overall accuracy (OA) and average accuracy (AA) under different labeling ratios. The best OA and AA at each ratio (except 1%) and their average values are highlighted in bold.
@@
TABLE VI: Comparison results on the AID dataset among HAL, various AL strategies, SSL methods, and their integrated variants (HSSAL). The table reports overall accuracy (OA) and average accuracy (AA) under different labeling ratios. The best OA and AA at each ratio (except 1%) and their average values are highlighted in bold.
@@
TABLE VII: Comparison results on the NWPU-RESISC45 dataset among HAL, various AL strategies, SSL methods, and their integrated variants (HSSAL). The table reports overall accuracy (OA) and average accuracy (AA) under different labeling ratios. The best OA and AA at each ratio (except 1%) and their average values are highlighted in bold.
@@
  • Entropy[16]: Queries samples with the highest predictive entropy, reflecting maximum model uncertainty.
  • Margin[66]: Selects samples with the smallest difference between the top two class probabilities, focusing on ambiguous predictions near the decision boundary.
  • BALD[18]: Estimates both epistemic and aleatoric uncertainty by maximizing the mutual information between predictions and model parameters. We adopt Monte Carlo dropout with $T=50$ stochastic forward passes on the last layer to approximate the acquisition score.
  • CoreSet[19]: Selects representative samples via $k$-center clustering in the feature space. The image features are the last-layer [CLS] tokens from the DINOv2-Small encoder, identical to those used in HAL.
  • BADGE[20]: Combines uncertainty and diversity by clustering gradient embeddings with $k$-means++. It operates on the same DINOv2 [CLS] features as CoreSet and HAL.
@@
  • FixMatch[13]: A representative weak-to-strong consistency method that generates pseudo-labels from weakly augmented images and enforces consistency on their strongly augmented counterparts when confidence exceeds a fixed threshold $\tau=0.95$.
  • FlexMatch[14]: Builds on FixMatch by introducing a class-wise adaptive threshold $\tau_{c}$ according to the learning progress of each class to reduce confirmation bias.
  • FreeMatch[15]: Further improves FixMatch by dynamically adjusting both global and class-specific thresholds based on prediction history for adaptive pseudo-labeling.
  • SoftMatch[30]: Extends the FixMatch paradigm by replacing the hard threshold with Gaussian confidence weighting, enabling smoother exploitation of uncertain samples.
As shown in Tables VVII, among all AL baselines, the proposed HAL consistently achieves the highest OA and AA across different labeling ratios. The close alignment between OA and AA indicates that HAL not only selects the most informative samples but also maintains class balance during the annotation process. This advantage mainly stems from its hierarchical sample selection mechanism, which ensures both representativeness and diversity in sample selection from the unlabeled pool.